There are 26 chocolates in a box, all identically shaped. There are 7 filled with nuts, 9 with caramel, and 10 are solid chocolate. You randomly select one piece, eat it, and then select a second piece. Find the probability of selecting 2 solid chocolates in a row?
Answer by checkley77(12569)

10/26*9/25=90/650=.13846 or 13.846% chance of getting 2 solid chocolate bars.